# Learned Kernels for Sparse, Interpretable, and Efficient Medical Time Series Processing

### v3

@@ 8 Code Availability
The code used for these experiments can be found at the following publically available GitHub repository: https://github.com/SullyChen/SMoLK.
9 Data Availability
We refer to the original works [46, 47, 48, 49, 51, 50] for the unprocessed data, as well as maintain a copy of the pre-processed data in our GitHub repository mentioned above.
10 Inclusion and Ethics
@@
Appendix A Baseline Methods
A.1 Convolutional Sliding Window
A.2 Template-Matching
We follow the same approach as Guo et al. [1], which was inspired by Lim et al.[54]. Signals were first divided from the PPG-DaLiA training set into separate pulses via peak detection. 10 clean (artifact-free) pulses were then selected to serve as our standard templates for comparison with test pulses. Each pulse from the test data was compared to these 10 templates, using the dynamic time warping (DTW) distance metric to measure similarity. The smallest DTW distance (out of the 10 comparisons with the templates) was then identified. The range of the DTW distance function is $[0,\infty)$, but empirically we find that the DTW distances between our templates and the PPG signals fall mostly within the range $[0,10]$. We created a binary predicted label by labeling a segment as an artifact if the minimum DTW distance is at least 1. Other samples were given “non-artifact” predicted labels. The threshold of DTW was chosen by testing thresholds between 0 and 10 over the entire train set and choosing the threshold that produced the highest DICE score (which was 1). To summarize, if the minimum DTW distance $\alpha$ exceeds our threshold, we classify the entire pulse (and all its timesteps) as an artifact.
A.3 GradCAM & SHAP
In this baseline experiment, the Resnet-34 architecture proposed by Dai et al (2016) was employed for 1D binary classification (identifying ‘clean’ or ‘artifact’). This architecture, traditionally used for image classification, was repurposed for time series analysis via GradCAM [42]. Considering the relatively small size of the training dataset, transfer learning was performed on a pre-trained Resnet34-1D PPG signal quality classifier, as described by Zhang et al (2021) [55]. This pre-trained model was originally trained on the UCSF PPG dataset according to Pereira et al (2019) [56]. The last two residual blocks, global average pooling, and the final dense layer were subsequently retrained by Guo et al. [1].
To train a classification model, ground truth labels were created such that any signal that contained artifact timesteps was labeled as an artifact, while all other signals were labeled as non-artifacts. After this labeling process, the training set consisted of 175 clean signals and 3261 artifact signals.
While the Resnet34-1D network is fundamentally a classifier, our goal is instead to obtain segmentation labels. Assuming that the model would focus on artifacts when tasked with predicting an artifact signal, we employed two post-hoc explanatory methods to generate segmentation masks: SHAP [43] and GradCAM [42].
A.4 Segade
A.5 Tiny-PPG
A.6 ResNet-16
Appendix B Post-Processing Filters
@@ Appendix C Data Limitations
Appendix D Pruning Results
@@
Appendix E Robustness to Noise
To test our model’s robustness to noise, we slowly add increasing amounts of Gaussian noise to our PPG segmentation test sets and evaluate the difference between the performance on the noisy signal and the original signal 9. We find that our learned kernel models are robust to large amounts of Gaussian noise, showing little degradation with a noticeably noisy signal.

### v4

@@ 8 Code Availability
The code has been made publicly available at https://doi.org/10.5281/zenodo.13117608, our GitHub repository [64]: https://github.com/SullyChen/SMoLK.
9 Acknowledgements
The authors would like to thank Anish Karpurapu for helpful discussions surrounding the interpretation of electrocardiograms. This work was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health under grant number R01HL166233 (XH). We gratefully acknowledge this support.
10 Competing Interests Statement
The authors declare no competing interests.
11 Author Contributions Statement
@@
Appendix A Post-Processing Filters
@@
Appendix B Scaling Laws
We explore the scaling behavior of our approach by gradually increasing the number of kernels and evaluating the performance on the test set. Our main objective is to understand the relationship between the parameter count and test-set performance, which helps us identify the potential limits and robustness of our approach.
We note that it is possible that a theoretical performance limit exists due to mislabeling in the test set; this is discussed further in the following section.
@@ Appendix C Data Limitations
Appendix D Robustness to Noise
To test our model’s robustness to noise, we slowly add increasing amounts of Gaussian noise to our PPG segmentation test sets and evaluate the difference between the performance on the noisy signal and the original signal (Figure 9). We find that our learned kernel models are robust to large amounts of Gaussian noise, showing little degradation with a noticeably noisy signal.
Appendix E Pruning Results